# Causality Pursuit from Heterogeneous Environments via Neural Adversarial Invariance Learning

### v1

@@ Abstract
Statistics suffers from a fundamental problem, “the curse of endogeneity” – the regression function, or more broadly the prediction risk minimizer with infinite data, may not be the target we wish to pursue. This is because when complex data are collected from multiple sources, the biases deviated from the interested (typically causal) association inherited in individuals or sub-populations are not expected to be canceled. Traditional remedies are of hindsight style and restrictive in being tailored to prior knowledge like untestable cause-effect structures, resulting in methods that risk severe model misspecification and lack scalable applicability. This paper seeks to offer a purely data-driven and universally applicable method that only uses the heterogeneity of the biases in the data themselves rather than following pre-offered and task-specific commandments. Such an idea is formulated as a nonparametric invariance pursuit problem, whose goal is to unveil the invariant conditional expectation $m^{\star}(x)\equiv\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}=x_{S^{\star}}]$ with unknown important variable set $S^{\star}$ across heterogeneous environments $e\in\mathcal{E}$. Under the structural causal model framework, $m^{\star}$ can be interpreted as pragmatic causality in general and it will coincide with the exact causal mechanism when there is enough heterogeneity. The paper contributes to proposing a novel framework, Focused Adversarial Invariance Regularized (FAIR) Estimation, formulated as a single minimax optimization program that can solve the general invariance pursuit problem. The adversary model in the framework serves the role of driving the predictor model towards a prediction-invariant solution adaptively. As illustrated by the unified non-asymptotic analysis, our adversarial estimation framework can attain provable sample-efficient estimation akin to standard regression under a minimal identification condition for various tasks and models. As an application, the Focused Adversarial Invariance Regularized Neural Network (FAIR-NN) estimator realized by two neural network classes is highlighted as the first approach to attain statistically efficient estimation in general nonparametric invariance learning.
@@ 1 Introduction
A fundamental problem in statistics and machine learning is to predict the response variable $Y$ based on explanatory covariates denoted as $X\in\mathbb{R}^{d}$. The objective often centers on estimating the regression function $m_{0}(x)=\mathbb{E}[Y|X=x]$, which minimizes the population $L_{2}$ risk
$\mathsf{R}(m)=\int|y-m(x)|^{2}\mu_{0}(dx,dy).$
based on an i.i.d. sample $\{(X_{i},Y_{i})\}_{i=1}^{n}$ drawn from $\mu_{0}$. Tracing back to pioneering works in the early nineteenth century (Legendre,, 1805; Gauss,, 1809), numerous statistical regression techniques have been developed to estimate the regression function $m_{0}$ from real-world data. By carefully balancing the bias and variance, the fitted model not only offers decent predictions for unseen data that is drawn from $\mu\approx\mu_{0}$ but also furnishes insights into the underlying mechanism between objects grounded in the real world. The above model is referred to as nonparametric regression (Fan & Gijbels,, 1996; Györfi et al.,, 2002; Tsybakov,, 2009) when it is agnostic about the functional form of $m$.
Climbing toward the hill of the efficient estimation of $m_{0}$ is challenging. When there are no additional low-dimension structures, estimating $m_{0}$ would suffer from the “curse of dimensionality” (Stone,, 1982). Unfortunately, the function structure may not often be known to us, whereas, the emergence of algorithmic (Breiman,, 2001) nonparametric techniques, e.g., neural networks, in the age of data make it possible to adaptively circumvent the curse of dimensionality. Compared with previous structural methods like additive models (Stone,, 1985) that pre-assume some low-dimensional structures, the neural network can obtain a comparable estimation error and circumvent the “curse of dimensionality” without the knowledge of the specific low-dimensional function structure ahead (Bauer & Kohler,, 2019; Schmidt-Hieber,, 2020; Fan & Gu,, 2022), and will not suffer from model misspecification. Such an algorithmic nature provides neural network estimators with scalable applicability in various prediction tasks and data structures.
@@
Though the progress we made in climbing towards $m_{0}$ is breathtaking, a more fundamental question is – are we climbing the right hill? To be specific, is $m_{0}$ always the target we wish to pursue, especially when gathering extensive variables for fitting or potentially collecting heterogeneous data from multiple sources? In numerous scenarios, $m_{0}$ itself may offer unstable predictions on unseen data from distributions where $\mu\neq\mu_{0}$, and risks misattributions and possible false scientific conclusions. Consider the simple thought experiment where we aim to classify an object in a picture as either a cow $(Y=1)$ or a camel $(Y=0)$ using two provided features $X_{1}$ (body shape) and $X_{2}$ (background color). Given our dataset drawn from the distribution $\mu_{0}$, the cows usually appear on green grass, while camels often stay on yellow sand. Consequently, the conditional expectation $m_{0}(x_{1},x_{2})=\mathbb{E}_{\mu_{0}}[Y|X_{1}=x_{1},X_{2}=x_{2}]$ might be heavily dependent on $x_{2}$. Such a model is problematic both for prediction and attribution. Its application in a setting with a different background such as zoos would lead to unreliable predictions. Furthermore, attributing the determination of an object to the background that surrounds it also contradicts our understanding of causality. In the above case, we may prefer $m_{\star}(x)=\mathbb{E}[Y|X_{1}=x_{1}]$ for prediction and attribution as we know the underlying causal mechanisms.
@@
We refer to the above problem as the “curse of endogeneity” in that the conditional expectation of the residual for the “potential” interested $m_{\star}$ is not zero given all the explanatory variables, i.e., $\mathbb{E}[Y-m_{\star}|X]\neq 0$, leading to a misalignment between $m_{0}$ and $m_{\star}$, i.e., $m_{0}(X)-m_{\star}(X)\neq 0$. Hence traditional nonparametric regression and machine learning for estimating $m_{0}$ will result in an unsatisfactory solution.
@@
Existing methods to overcome the problem are mainly restricted to causal inference methods, in which some “prior” structural assumptions or a causal graph are often pre-offered before the inference. Such prior knowledge can be formally encoded in estimation within the potential outcome (Rubin,, 1974), or structural causal model (SCM) (Glymour et al.,, 2016) framework. The key idea is that for a specific causal estimation task, one can represent some causal estimand through a statistical estimand given the prior knowledge, i.e., identification conditions. The statistical estimand can be subsequently estimated using regression techniques. Examples include the average treatment effect estimation (Robins et al.,, 1994) under the unconfoundedness condition, and instrument variable regression (Reiersøl,, 1945; Heckman,, 2008) under the valid instrument condition. It is worth pointing out that all of these methods are tailored to prior task-specific and untestable knowledge like cause-effect relationships between the variables, and thus can be categorized as the “structural methods” as a remedy of the “curse of endogeneity”. Specifically, the “causality skeleton” is fully shaped by the prior commandments in the first stage, while the collected data and the statistical estimation methods are only responsible for filling the “association flesh” in the second stage. This limitation prevents the conclusion they drawn from going beyond hindsight in general and hinders them from scalable applicability because it requires domain expert knowledge for very complicated data, and risks more serious model misspecification since such knowledge is not testable and hence impossible to be falsified in general.
In this paper, we consider shaping both the “causality skeleton” and “association flesh” from the data themselves. The goal is to answer the following fundamental question:
@@
Can we design methods that can circumvent the “curse of endogeneity” without thesupervision of cause-effect $\begin{split}&\text{\it Can we design methods that can circumventthe ``curse of endogeneity'' without the}\\ &\text{\it supervision of cause-effect knowledge?}\end{split}$ (Q)
Without the supervision of prior causal structural knowledge, we leverage the principle of how humans understand causality: the causal association is the one that consistently occurs in the past, now, and (potentially) in the future, or more broadly, in diverse environments. The main contribution of this paper is to propose a unified, algorithmic methodological framework to circumvent the “curse of endogeneity” in a purely data-driven way. This algorithmic nature provides it with scalable applicability and is free of model misspecification in a similar manner to neural networks as the remedy of the “curse of dimensionality”.
@@ 1.1 The Canonical Model under Study
Let us revisit the thought experiment from the perspective of a hyper-intelligent alien, Alice. Alice knows nothing about cows and camels except for 1000 images with annotated labels highly associated with the background, for example, $r=90\%$ cows/camels on grass/sand. It’s impossible for her to claim that the background cannot determine the object given this limited information. However, if she receives another set of 1000 images, where $r=70\%$ cows/camels on grass/sand, she might begin to question the determinative role of the background. Her skepticism will intensify when more and more datasets with varying $r$ are offered.
When there is no supervision of the cause-effect relationship, the observation from heterogeneous sources is essential in the nonparametric setting. In this context, this paper considers the following multi-environment regression model. Let $\mathcal{E}$ be the set of sources/environments. For each environment $e\in\mathcal{E}$, we observe $n$ [^1] We point out that the same $n$ is just for expository simplicity, the extension to varying $n^{(e)}$ is straightforward. i.i.d. pairs $(X_{1}^{(e)},Y_{1}^{(e)}),\ldots,(X_{n}^{(e)},Y_{n}^{(e)})\sim\mu^{(e)}$, where $\mu^{(e)}$, the joint distribution of $(X^{(e)},Y^{(e)})$, satisfies
@@
Here $S^{\star}$, the unknown true important variable set, and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$, the target regression function, are both invariant across different environments; the joint distributions $\mu^{(e)}$ can vary. We aim to estimate the invariant regression function $m^{\star}$ using data $\{\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\}_{e\in\mathcal{E}}$ from $|\mathcal{E}|$ heterogeneous environments. We refer to the above problem as nonparametric invariance pursuit.
@@
Here we temporarily refrain from causal discussions and frame it as a pure statistical estimation problem. Under particular scenarios, such a problem can be instantiated to causal discovery in the SCM framework (Peters et al.,, 2016), and transfer learning with a more realistic assumption (Rojas-Carulla et al.,, 2018); see the details in Section A.1. We offer a rigorous and comprehensive interpretation of what $S^{\star}$ is in the SCM with interventions on $X$ in Section 4.2. It is also notable to mention that model (1.1) only requires invariance in the first moment instead of full distributional invariance, i.e., $\varepsilon^{(e)}\sim F_{\varepsilon}\perp\!\!\!\perp X_{S^{\star}}^{(e)}$, as typically required for causal discovery (Peters et al.,, 2016). The invariance structure in (1.1) is more realistic in that it allows for between-environment heteroscedastic errors. The relaxation we adopted broadens the scope for various applications, including binary classification.
@@
It is important to note that the objective of standard nonparametric regression generally diverges from our target $m^{\star}$, i.e., $\mathbb{E}[Y^{(e)}|X^{(e)}=x]\neq m^{\star}(x_{S^{\star}})$. This discrepancy arises because the exogeneity condition $\mathbb{E}[\varepsilon^{(e)}|X_{S^{\star}}^{(e)}]=0$ only applies to the true important variables, while $\mathbb{E}[\varepsilon^{(e)}|X^{(e)}]\neq 0$. Such a “curse of endogeneity” problem is the main challenge we need to address. To identify $m^{\star}$, we exploit the invariance structure
$\forall e,e^{\prime}\in\mathcal{E}\qquad\mathbb{E}[Y^{(e)}|X_{S^{\star}}^{(e)}]\equiv\mathbb{E}[Y^{(e^{\prime})}|X_{S^{\star}}^{(e^{\prime})}].$ (1.2)
1.2 Contributions of the Paper
@@
In this paper, we propose a unified estimation framework – the Focused Adversarial Invariance Regularized (FAIR) estimator. This estimator regularizes the user-specified risk loss $\ell(y,v)$ using a novel regularizer. Specifically, the FAIR estimator is the solution of the following minimax optimization program
@@
$\min_{g\in\mathcal{G}}\max_{f^{(e)}\in\mathcal{F}_{S_{g}},\forall e\in\mathcal{E}}\underbrace{\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\ell(Y,g(X))\right]}_{\mathsf{R}(g)}+\gamma\underbrace{\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\{Y-g(X)\}f^{(e)}(X)-\{f^{(e)}(X)\}^{2}/2\right]}_{\mathsf{J}(g,\{f^{(e)}\}_{e\in\mathcal{E}})}.$
@@
Here $\ell(\cdot,\cdot)$ is a loss whose population solution leads to the conditional expectation, $\gamma>0$ is the regularization hyper-parameter to be determined, $(\mathcal{G},\mathcal{F})$ are the function classes to be specified by the user satisfying $\mathcal{G}\subseteq\mathcal{F}$, $\mathcal{F}_{S_{g}}=\{f\in\mathcal{F}:f(x)=h(x_{S_{g}})\text{ for some }h:\mathbb{R}^{|S_{g}|}\to\mathbb{R}\}$, and $S_{g}=\mathrm{supp}(g)$ represents the index set of the variables $g$ depends on. The function class $\mathcal{F}$ serves as the testing function class for the prediction functions in $\mathcal{G}$. From a high-level viewpoint, solving the minimax optimization program of $\mathsf{J}$ promotes the invariance in a similar spirit to (1.1), while minimizing $\mathsf{R}$ with respect to $g$ prevents it from collapsing to conservative solutions. For the empirical counterpart, we solve a similar minimax optimization program that substitutes $\mathbb{E}_{\mu^{(e)}}[\cdot]$ with the corresponding sample means. We highlight our core contributions as follows.
At the technical level, we propose a new method that can discover the invariant regression function, i.e. to solve a generalized version of the problem in Section 1.1, in an algorithmic manner. Specifically, the newly proposed FAIR estimator can be obtained by simply solving a minimax optimization program blind to the cause-effect knowledge. The method is simple – it is just one optimization objective complemented by one extra hyper-parameter $\gamma$. Moreover, it is universal – it accommodates any loss $\ell(\cdot,\cdot)$ whose unique optimal solution aligns with the conditional expectation, and can be seamlessly integrated by a variety of machine learning algorithms. Therefore, our method, which can also be regarded as a general estimation framework, homogenizes various tasks such as regression and classification, and diverse models including neural networks (Barron,, 1994) and the spline-based MARS algorithm (Friedman,, 1991). The framework encapsulates previously studied methods as particular instances. For example, with both $\mathcal{F}$ and $\mathcal{G}$ as linear function classes, it turns into a stabilized variant of the EILLS estimator (Fan et al.,, 2023). When $|\mathcal{E}|=1$, it is similar to the FGMM estimator (Fan & Liao,, 2014) under a certain choice of $(\mathcal{G},\mathcal{F})$. We also conduct a unified non-asymptotic theoretical analysis for our method of estimation. Our main result extends the theoretical findings in Fan et al., (2023) to a broad spectrum of $(\mathcal{G},\mathcal{F},\ell)$ specifications.

### v2

@@ Abstract
Pursuing causality from data is a fundamental problem in scientific discovery, treatment intervention, and transfer learning. This paper introduces a novel algorithmic method for addressing nonparametric invariance and causality learning in regression models across multiple environments, where the joint distribution of response variables and covariates varies, but the conditional expectations of outcome given an unknown set of quasi-causal variables are invariant. The challenge of finding such an unknown set of quasi-causal or invariant variables is compounded by the presence of endogenous variables that have heterogeneous effects across different environments, including even one of them in the regression would make the estimation inconsistent. The proposed Focused Adversial Invariant Regularization (FAIR) framework utilizes an innovative minimax optimization approach that breaks down the barriers, driving regression models toward prediction-invariant solutions through adversarial testing. Leveraging the representation power of neural networks, FAIR neural networks (FAIR-NN) are introduced for causality pursuit. It is shown that FAIR-NN can find the invariant variables and quasi-causal variables under a minimal identification condition and that the resulting procedure is adaptive to low-dimensional composition structures in a non-asymptotic analysis. Under a structural causal model, variables identified by FAIR-NN represent pragmatic causality and provably align with exact causal mechanisms under conditions of sufficient heterogeneity. Computationally, FAIR-NN employs a novel Gumbel approximation with decreased temperature and stochastic gradient descent ascent algorithm. The procedures are convincingly demonstrated using simulated and real-data examples.
@@ 1 Introduction
A fundamental problem in statistics and machine learning is to predict the response variable $Y$ based on explanatory covariates denoted as $X\in\mathbb{R}^{d}$ using collected data. The objective often centers on estimating the regression function $m_{0}(x)=\mathbb{E}[Y|X=x]$, which minimizes the population $L_{2}$ risk $\mathsf{R}(m)=\int|y-m(x)|^{2}\mu_{0}(dx,dy)$, starting from the pioneering work of least squares in Legendre, (1805); Gauss, (1809). In the age of data, the problem of achieving sample-efficient estimation of $m_{0}$ was extensively studied. There are a lot of structural methods attempting to exploit the low-dimensional structure such as sparsity, low-rankness and additivity, and design corresponding optimal methods tailored to that assumed structure (Hastie et al.,, 2009; Wainwright,, 2019; Fan et al.,, 2020). However, these methods lack scalable applicability and suffer from model misspecification due to their reliance on imposed structures. As an alternative, algorithmic methods (Breiman,, 2001) like neural networks can be adaptive to the low-dimensional structure efficiently (Schmidt-Hieber,, 2020; Fan & Gu,, 2024) with no supervision of function structure. This nature endows them with universal applicability across various tasks and data.
@@
Despite many celebrated efforts in the efficient estimation of $m_{0}$ or its variants like quantile function, the ultimate goal is to utilize observations to fit a model capable of making decent predictions on unseen data, elucidating the causal relationships among variables, and guiding decision-making in real-world scenarios. We instinctively regard $m_{0}$ as such a target function for achieving decent prediction and causal attribution. However, this can be flawed: $m_{0}$ can produce unstable predictions on unseen data and risk false scientific conclusions in numerous cases. Consider a simple thought experiment where we aim to classify an object in a picture as either a cow $(Y=1)$ or a camel $(Y=0)$ using two provided features $X_{1}$ (body shape) and $X_{2}$ (background color). In the data we collected from $\mu_{0}$, the cows usually appear on green grass, while camels often stay on yellow sand. Consequently, the conditional expectation $m_{0}(x_{1},x_{2})=\mathbb{E}_{\mu_{0}}[Y|X_{1}=x_{1},X_{2}=x_{2}]$ would be heavily dependent on $x_{2}$. Such a model is problematic both for prediction and attribution. Its application in a setting with a different background such as zoos would lead to unreliable predictions. Furthermore, attributing the determination of an object to the background surrounding it also contradicts our understanding of causality. In the above case, we may prefer $m_{\star}(x)=\mathbb{E}[Y|X_{1}=x_{1}]$ for prediction and attribution as we know the causal mechanisms.
@@
We refer to the above problem as the “curse of endogeneity” in that the conditional expectation of the residual for the “potential” interested (causal) $m_{\star}$ is not zero given all the explanatory variables, i.e., $\mathbb{E}[Y-m_{\star}(X)|X]\neq 0$, leading to a misalignment between $m_{0}$ and $m_{\star}$, i.e., $m_{0}(X)-m_{\star}(X)\neq 0$. Hence traditional regression techniques for estimating $m_{0}$ will result in an unsatisfactory solution.
@@
Causal inference methods offer structural remedies to the curse of endogeneity. These methods are structural in that they are tailored to pre-assumed, task-specific, and untestable identification conditions or causal-effect knowledge. This prior knowledge can be formally encoded in the potential outcome (Rubin,, 1974), or structural causal model (Glymour et al.,, 2016) framework and fully shaped the “causality skeleton” that exactly determines the causal estimand of interests by some statistical estimand, and the “association flesh” of the latter can be further estimated via structural or algorithmic regression techniques. Examples include estimating the average treatment effect (Robins et al.,, 1994) and conditional average treatment effects (Athey et al.,, 2019; Kennedy et al.,, 2024) under the unconfoundedness condition. These methods’ reliance on prior knowledge limits their scalable use, exposes them to severe model misspecification, and prevents their drawn conclusion from going beyond hindsight because it is impossible to falsify (Popper,, 2005) these assumptions using data.
This paper aims to answer the following fundamental question:
@@
Can we design methods that can algorithmically circumvent the “curse of endogeneity”without the supervision of cause-effect $\begin{split}&\text{\it Can we design methods that can algorithmically circumvent the ``curse of endogeneity''}\\ &\text{\it without the supervision of cause-effect knowledge?}\end{split}$ (Q)
Without prior causal structural knowledge, we leverage the principle of how humans understand causality: the causal association consistently occurs in the past, now, and (potentially) future, or more broadly, in diverse environments. In other words, we pursue certain data-driven or data-shaped causality that is invariant across diverse environments, this is essentially what one can pursue based only on observed data without prior knowledge. Hence we do not differentiate the concepts of invariance and (data-driven) causality in this paper. Levering the invariance principle, we propose a unified and algorithmic framework for causality pursuit that is robust to model misspecification based on data from multiple environments. Though the proposed data-driven causality is conceptually different from previous knowledge-based causality that pre-assumes the ground truth, these two types of causality can coincide when the heterogeneity of environments is sufficient.
@@ 1.1 The Canonical Model under Study
Let us revisit the thought experiment from the perspective of a hyper-intelligent alien, Alice. Alice knows nothing about cows and camels except for 1000 images with annotated labels highly associated with the background, for example, $r=90\%$ cows/camels on grass/sand. It’s impossible for her to know that the background cannot determine the object given this limited information. In other words, both $X_{1}$ and $X_{2}$ can be regarded as causality out of pragmatic considerations. However, if she receives another set of 1000 labeled images, where $r=70\%$ cows/camels on grass/sand, she might begin to question the causality role of the background: the emerging evidence of the varying associations between $X_{2}$ and $Y$ falsify the hypothesis that $X_{2}$ is causality if she believes that causality persists across diverse environments.
When there is no supervision of the cause-effect relationship, the observation from heterogeneous sources is essential. We consider the following multi-environment regression problem that mimics human causality learning. Let $\mathcal{E}$ be the set of sources/environments. For each environment $e\in\mathcal{E}$, we observe $n$ i.i.d. data $(X_{1}^{(e)},Y_{1}^{(e)}),\ldots,(X_{n}^{(e)},Y_{n}^{(e)})\sim\mu^{(e)}$, where $\mu^{(e)}$, the joint distribution of $(X^{(e)},Y^{(e)})$, satisfies
@@
Here $S^{\star}$, the unknown true important variable set, and $m^{\star}:\mathbb{R}^{|S^{\star}|}\to\mathbb{R}$, the target regression function, are both invariant across different environments; but the joint distributions $\mu^{(e)}$ can vary. We aim to learn the set of quasi-causal variables $S^{\star}$ and estimate the invariant regression function $m^{\star}$ using data $\{\{(X_{i}^{(e)},Y_{i}^{(e)})\}_{i=1}^{n}\}_{e\in\mathcal{E}}$ from $|\mathcal{E}|$ heterogeneous environments. The same $n$ in the problem formulation is just for expository simplicity, the extension to varying $n^{(e)}$ is straightforward. We refer to the above problem as nonparametric invariance pursuit or nonparametric causality pursuit exchangeably, as based on the data alone, without prior knowledge, we can not differentiate these concepts.
@@
Here, we temporarily refrain from causal discussions. Under particular scenarios, such a problem can be instantiated to causal discovery in the Structural Causal Model (SCM) framework (Peters et al.,, 2016) and transfer learning with a more realistic assumption (Rojas-Carulla et al.,, 2018); see the details in Section A.1. We offer in Section 3 a rigorous and comprehensive interpretation of what $S^{\star}$ is in the SCM with interventions on $X$. It is also notable to mention that model (1.1) only requires invariance in the first moment instead of full distributional invariance, i.e., $\varepsilon^{(e)}\sim F_{\varepsilon}$ and independent of $X_{S^{\star}}^{(e)}$, as typically required for causal discovery (Peters et al.,, 2016). It is more realistic and allows for between-environment heteroscedastic errors.
@@
It is important to note that the standard nonparametric regression generally diverges from our target $m^{\star}$, i.e., $\mathbb{E}[Y^{(e)}|X^{(e)}=x]\neq m^{\star}(x_{S^{\star}})$. This discrepancy arises because $\mathbb{E}[\varepsilon^{(e)}|X^{(e)}]\neq 0$. Such a “curse of endogeneity” problem is the main challenge we need to address. Including even one of endogenous spurious variables, for example, $X_{2}$ background color in the above thought experiment, in the regression function will create an inconsistent estimation of $m^{\star}$. Thus, it is essential to design an algorithm to eliminate all endogenous spurious variables.
1.2 Our Algorithmic Remedy: FAIR Estimation
@@
This paper proposes a unified estimation framework – the Focused Adversarial Invariance Regularized (FAIR) estimator. It regularizes the user-specified risk loss $\ell(y,v)$ by a novel regularizer. Specifically, the FAIR estimator is the solution of the following minimax optimization program
@@
$\min_{g\in\mathcal{G}}\max_{f^{(e)}\in\mathcal{F}_{S_{g}},\forall e\in\mathcal{E}}\underbrace{\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\ell(Y,g(X))\right]}_{\mathsf{R}(g)}+\gamma\underbrace{\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\{Y-g(X)\}f^{(e)}(X)-\{f^{(e)}(X)\}^{2}/2\right]}_{\mathsf{J}(g,\{f^{(e)}\}_{e\in\mathcal{E}})}.$ (1.2)
@@
Here $\ell(\cdot,\cdot)$ is a loss whose population solution leads to the conditional expectation, $\gamma>0$ is the regularization hyper-parameter to be determined, $(\mathcal{G},\mathcal{F})$ are the function classes to be specified by the user satisfying $\mathcal{G}\subseteq\mathcal{F}$. The first part is the risk minimization, and the second component is the test of exogeneity of the variables $S_{g}=\mathrm{supp}(g)$ used by the regression function $g$, where $\mathcal{F}_{S_{g}}=\{f\in\mathcal{F}:f(x)=h(x_{S_{g}})\text{ for some }h:\mathbb{R}^{|S_{g}|}\to\mathbb{R}\}$ is the testing function class for the prediction functions in $\mathcal{G}$ that only “focuses” on the variables $S_{g}$ that $g$ used. Two useful classes of functions are linear and square-integrable classes for $(\mathcal{G},\mathcal{F})$, which correspond respectively to linear models and nonparametric regression models; see Section 4.1 for additional details. Note that the second component is nonnegative after maximization by comparing with $f^{(e)}=0$ so that the penalty is nonnegative. For the empirical counterpart, we solve a similar minimax optimization program that substitutes $\mathbb{E}_{\mu^{(e)}}[\cdot]$ with the corresponding sample means.
To see why such a FAIR penalty works, let us consider the nonparametric regression setting in which =$\mathcal{F}=\{f:\mathbb{E}_{\mu^{(e)}}[f^{2}(X_{S_{g}})]<\infty\}$. By conditioning on $X_{S_{g}}$, for $f^{(e)}\in\mathcal{F}_{S_{g}}$, we have
$\mathbb{E}_{\mu^{(e)}}\left[\{Y-g(X)\}f^{(e)}(X)\right]=\mathbb{E}_{\mu^{(e)}}\left[\left\{\mathbb{E}_{\mu^{(e)}}[Y|X_{S_{g}}]-g(X)\right\}f^{(e)}(X)\right].$
Then, the supremum in (1.2) can be explicitly found and the objective now becomes
$\min_{g\in\mathcal{G}}\mathsf{R}(g)+\gamma\cdot\mathsf{J}^{\star}(g)\qquad\text{with}\qquad\mathsf{J}^{\star}(g)=\frac{1}{2}\sum_{e\in\mathcal{E}}\mathbb{E}_{\mu^{(e)}}\left[\left|g(X)-\mathbb{E}_{\mu^{(e)}}[Y|X_{S_{g}}]\right|^{2}\right].$ (1.3)
Therefore, $g(X)=m^{\star}(X_{S^{\star}})$ is a minimax solution.
To motivate (1.2), let us first consider the additional constraint $\mathbb{E}_{\mu^{(e)}}[f^{(e)}(X_{S_{g}})^{2}]=1$ so that the first part of the second component in (1.2) is basically the maximal correlation between the residual $\{Y-g(X_{S_{g}})\}$ and testing functions $f^{(e)}(X_{S_{g}})$. Hence, the criterion (1.2) is to find a set of variables $X_{S_{g}}$ as exogenous (weakly correlated) with the residuals as possible for all testing functions in $\mathcal{F}_{S_{g}}$. By the Lagrange multiplier method, the constrained maximization problem can be written as
$\max_{f^{(e)}\in\mathcal{F}_{S_{g}}}\mathbb{E}_{\mu^{(e)}}\left[\{Y-g(X)\}f^{(e)}(X)-\lambda\{f^{(e)}(X)\}^{2}\right].$
Choosing the multiplier $\lambda=1/2$ (justified in the above paragraph) gives rise to the object function (1.2).
FAIR penalty screens out all endogenous spurious variables when $\gamma$ is sufficiently large. This is easily seen when the penalty in (1.2) is not zero, such a $g$ is dominated by $g=m^{\star}$ when $\gamma$ is sufficiently large. After endogenous spurious variables, we can apply the commonly-used statistical variable selection methods (Hastie et al.,, 2009; Wainwright,, 2019; Fan et al.,, 2020) to further eliminate exogenous spurious or weak causal variables. In addition, we will show that under the SCM with arbitrary and nondegenerate interventions on $X$, our proposed FAIR estimator can unveil $S^{\star}$ being precisely expressed by the graph structure of the SCM, which can be interpreted as the “pragmatic” direct cause of the response $Y$ in general and will coincide with the direct causes if all the root children are intervened. The obtained result is clearly distinguished from what least squares, or even its worst-case variants like distribution robust optimization (Duchi & Namkoong,, 2021), Maximin (Meinshausen & Bühlmann,, 2015), can obtain. Our method indeed learns certain data-driven causality, while others cannot go beyond learning associations.
1.3 New Contributions
We propose a unified, algorithmic, and sample-efficient methodological framework that can discover the invariant regression function, i.e. to solve a generalized version of the problem in Section 1.1. The method is simple, universal, fully algorithmic, and sample-efficient: It is just one optimization objective (1.2) complemented by one extra hyper-parameter $\gamma$; it accommodates many losses and can be seamlessly integrated by various machine learning algorithms; it does not require any prior structural knowledge, and it is almost as statistically efficient as standard regression under various cases.